OpenStudy (anonymous):

Help please? Problem: 3√45^3 + 2√12 + √27-3√20x^3 Answers: a.)3x√5x+7√3 b.)3x√10x+7√6 c.)10x√8x d.)40x√x

jhonyy9 (jhonyy9):

3sqrt45^3 = 3sqrt(45*45*45) =3sqrt(45^2 *45) =3*45sqrt45 =135sqrt45 = =135sqrt(9*5) =135*3sqrt5 =405sqrt5 so can you continue the same from this ?

jhonyy9 (jhonyy9):

2sqrt12 =2*2sqrt3 =4sqrt3 sqrt27 =sqrt(9*3) =3sqrt3

OpenStudy (anonymous):

Thanks you guys! I'm reallly bad at this section. :(

jhonyy9 (jhonyy9):

3sqrt20x^3 =3sqrt(4*5*x^2 *x) =3*2*xsqrt5x =6xsqrt5x

jhonyy9 (jhonyy9):

405sqrt5 +4sqrt3 +3sqrt3 -6xsqrt5x = can you calcule this ?

jhonyy9 (jhonyy9):

so from these all result that 7sqrt3 need being in your right answer sure but if you check it my calcules with this great number 405sqrt5 --- so this dont appair there in options so i think that you have missed anything from your exercise